# Optimizing Parameters for Static Equilibrium of Discrete Elastic Rods with Active-Set Cholesky

### v1

@@ III-D Twisting
$E_{\mathrm{tw},i}(\mathbf{y}_{i})=\frac{1}{2}\left(\frac{s\bm{\gamma}_{i}\pi r^{4}}{(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)(\mathbf{m}_{i}-\mathbf{\bar{m}}_{i})^{2},$ (7)
with $\bm{\gamma}_{i}$ representing the twisting stiffness parameter, and $\mathbf{m}_{i}$ and $\mathbf{\bar{m}}_{i}$ the twist (including reference twist [7, 40]) and rest twist, respectively. The 11-dimensional gradient of (7) is
$\nabla E_{\mathrm{tw},i}(\mathbf{y}_{i})=\left(\frac{s\bm{\gamma}_{i}\pi r^{4}}{(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)(\mathbf{m}_{i}-\mathbf{\bar{m}}_{i})\nabla\mathbf{m}_{i}.$ (8)
IV Parameter Optimization
We simultaneously optimize parameters of the rest shape ($\mathbf{\bar{l}},\bm{\bar{\kappa}},\mathbf{\bar{m}}$) and material stiffness ($\bm{\alpha},\bm{\beta},\bm{\gamma}$) to achieve static equilibrium of an elastic strand. While one typically specifies stiffness coefficients directly as $\mathbf{c}_{\mathrm{st}}(=s\bm{\alpha})$, $\mathbf{c}_{\mathrm{be}}(=s\bm{\beta})$, and $\mathbf{c}_{\mathrm{tw}}(=s\bm{\gamma})$, we instead optimize $\bm{\alpha}$, $\bm{\beta}$, and $\bm{\gamma}$. This approach addresses the significant scale differences between the stiffness coefficients ($\mathbf{c}_{\mathrm{st}}$, $\mathbf{c}_{\mathrm{be}}$, $\mathbf{c}_{\mathrm{tw}}$ on the order of $10^{9}$ [7, 40]) and rest shape parameters ($\mathbf{\bar{l}},\bm{\bar{\kappa}},\mathbf{\bar{m}}$ up to 10), enabling stable convergence of numerical optimizers and mitigating numerical issues. To ensure similar value scales, we initialize the stiffness scaling constant $s$ as $s=\frac{1}{3(N-2)}\sum_{i=1}^{N-2}(\mathbf{c}_{\mathrm{st},i}+\mathbf{c}_{\mathrm{be},i}+\mathbf{c}_{\mathrm{tw},i})$, and subsequently set $\bm{\alpha}=\mathbf{c}_{\mathrm{st}}/s$, $\bm{\beta}=\mathbf{c}_{\mathrm{be}}/s$, and $\bm{\gamma}=\mathbf{c}_{\mathrm{tw}}/s$. While equivalent formulations could be derived through proper scaling of (9) and (10), our change of variables ensures consistent scales in the Jacobians (see Sec. IV-E), thus avoiding catastrophic cancellation and enhancing numerical stability.
Exploiting the one-dimensional structure and locally defined DER objectives of strands, we interleave the rest shape and stiffness parameters to ensure banded inner linear systems within Newton-type optimizers, similar to the approach used in forward simulations [7]. Additionally, we exclude $\mathbf{\bar{l}}_{0}$ on the fixed edge from the parameter optimization [5].
Furthermore, we optimize only two rest curvature DOFs per vertex $\bm{\bar{\kappa}}_{i,0}$ and $\bm{\bar{\kappa}}_{i,1}$ while excluding $\bm{\bar{\kappa}}_{i,2}$ and $\bm{\bar{\kappa}}_{i,3}$, in contrast to previous work [5]. This choice stems from the association of $\bm{\bar{\kappa}}_{i,0}$ and $\bm{\bar{\kappa}}_{i,2}$ with the second material frame, while $\bm{\bar{\kappa}}_{i,1}$ and $\bm{\bar{\kappa}}_{i,3}$ relate to the first [6]. Their overlapping force spaces lead to similar final values for $\bm{\bar{\kappa}}_{i,0}$ and $\bm{\bar{\kappa}}_{i,2}$ without expanding the force spaces necessary for static equilibrium. Including $\bm{\bar{\kappa}}_{i,2}$ and $\bm{\bar{\kappa}}_{i,3}$ would increase the optimization cost due to the additional DOFs. In practice, we synchronize these variables by setting $\bm{\bar{\kappa}}_{i,2}=\bm{\bar{\kappa}}_{i,0}$ and $\bm{\bar{\kappa}}_{i,3}=\bm{\bar{\kappa}}_{i,1}$ whenever $\bm{\bar{\kappa}}_{i,0}$ and $\bm{\bar{\kappa}}_{i,1}$ are updated. This approach effectively reduces the dimensionality of the rest curvature from 4D to 2D; however, we still use 4D rest curvatures in (5) to ensure the same dimensionality for curvature $\bm{\kappa}$ and rest curvature $\bm{\bar{\kappa}}$.
Concretely, we define the optimization variable for the parameters as $\mathbf{p}=(\bm{\bar{\kappa}}_{1,0},\bm{\bar{\kappa}}_{1,1},{\bm{\beta}}_{1},\mathbf{\bar{m}}_{1},{\bm{\gamma}}_{1},\mathbf{\bar{l}}_{1},{\bm{\alpha}}_{1},\ldots,\\ \bm{\bar{\kappa}}_{N-2,0},\bm{\bar{\kappa}}_{N-2,1},{\bm{\beta}}_{N-2},\mathbf{\bar{m}}_{N-2},{\bm{\gamma}}_{N-2},\mathbf{\bar{l}}_{N-2},{\bm{\alpha}}_{N-2})^{T}\in\mathbb{R}^{7N-14}$. Our central parameter optimization task is then formulated as a constrained minimization problem,
@@ IV Parameter Optimization
where $\mathbf{p}_{\mathrm{min}}$ and $\mathbf{p}_{\mathrm{max}}$ denote lower and upper bounds for $\mathbf{p}$ to ensure compliance with physical laws and limit significant parameter changes [5], and $\mathbf{c}(\mathbf{p})$ (defined in Sec. IV-B) is a nonlinear hard constraint to enforce zero net forces. We define $\mathbf{p}_{0}$ as the initial value for $\mathbf{p}$ before optimization, and $R(\mathbf{p})$ acts as a regularizer to minimize the deviation of $\mathbf{p}$ from $\mathbf{p}_{0}$ with a diagonal weight matrix $\mathbf{W}$. Below, we assume $\Delta t=1$ as it does not impact the optimization results [5].
IV-A Augmented Lagrangian Method
@@
To handle the nonlinear constraints $\mathbf{c}(\mathbf{p})=0$ in (9), we use ALM [15, 35] and define an augmented Lagrangian objective with a Lagrange multiplier $\bm{\lambda}$ and a penalty parameter $\rho$ as
@@
$L(\mathbf{p},\bm{\lambda})=R(\mathbf{p})+\bm{\lambda}^{T}\mathbf{c}(\mathbf{p})+\frac{\rho}{2}\left\lVert\mathbf{c}(\mathbf{p})\right\rVert_{2}^{2}.$ (10)
@@
$\mathbf{p},\bm{\lambda}=\operatorname*{arg\,min}_{\mathbf{p}_{\mathrm{min}}\leq\mathbf{p}\leq\mathbf{p}_{\mathrm{max}}}L(\mathbf{p},\bm{\lambda}),$ (11)
which we optimize by iteratively solving primal and dual subproblems [15]. The primal subproblem, with fixed $\bm{\lambda}^{k}$ for iteration index $k$ (initialized with $\bm{\lambda}^{0}=0$), is defined as
$\mathbf{p}^{k+1}=\operatorname*{arg\,min}_{\mathbf{p}_{\mathrm{min}}\leq\mathbf{p}\leq\mathbf{p}_{\mathrm{max}}}L(\mathbf{p},\bm{\lambda}^{k}),$ (12)
while the dual subproblem becomes a simple vector update:
$\bm{\lambda}^{k+1}=\bm{\lambda}^{k}+\rho\mathbf{c}(\mathbf{p}^{k+1}).$ (13)
@@
IV-B Zero Net Force Constraints
Given the total force due to the DER objectives, represented as $\mathbf{f}(\mathbf{p})=-\nabla E(\mathbf{q})$, it is natural to define $\mathbf{c}(\mathbf{p})=\mathbf{f}(\mathbf{p})$ to enforce zero net forces. However, in this case, the term $\frac{\rho}{2}\left\lVert\mathbf{c}(\mathbf{p})\right\rVert_{2}^{2}$ in (10) becomes equivalent to a quadratic penalty on the total force, which is numerically ill-conditioned and fails to accurately represent force contributions within the system [5]. Therefore, we define $\mathbf{c}(\mathbf{p})$ as
@@ IV-B Zero Net Force Constraints
Consequently, $L(\mathbf{p},\bm{\lambda})$ in (10) can be rewritten as
$L(\mathbf{p},\bm{\lambda})=R(\mathbf{p})+\bm{\lambda}^{T}\mathbf{M}^{-\frac{1}{2}}\mathbf{f}(\mathbf{p})+\frac{\rho}{2}\left\lVert\mathbf{f}(\mathbf{p})\right\rVert^{2}_{\mathbf{M}^{-1}}.$ (15)
[…]
IV-C Box Constraints
We impose box constraints on the optimization parameters to prevent stability problems and to more accurately control the resulting strand dynamics. Specifically, we define $\mathbf{\bar{l}}_{\mathrm{min}}\leq\mathbf{\bar{l}}\leq\mathbf{\bar{l}}_{\mathrm{max}}$ (except for $\mathbf{\bar{l}}_{0}$), $\bm{\bar{\kappa}}_{\mathrm{min}}\leq\bm{\bar{\kappa}}\leq\bm{\bar{\kappa}}_{\mathrm{max}}$, $\mathbf{\bar{m}}_{\mathrm{min}}\leq\mathbf{\bar{m}}\leq\mathbf{\bar{m}}_{\mathrm{max}}$, $\bm{\alpha}_{\mathrm{min}}\leq\bm{\alpha}\leq\bm{\alpha}_{\mathrm{max}}$, $\bm{\beta}_{\mathrm{min}}\leq\bm{\beta}\leq\bm{\beta}_{\mathrm{max}}$, and $\bm{\gamma}_{\mathrm{min}}\leq\bm{\gamma}\leq\bm{\gamma}_{\mathrm{max}}$. Unlike the penalty-based approach [5], which requires safety margins to mitigate the risk of violating physical laws (e.g., negative rest length and material parameters), we handle box constraints precisely using MPRGP [16]. Consequently, we set lower and upper bounds for $\mathbf{\bar{l}}$, $\bm{\alpha}$, $\bm{\beta}$, and $\bm{\gamma}$ as
$\mathbf{\bar{l}}_{\mathrm{min}}$ $=\epsilon,\quad\mathbf{\bar{l}}_{\mathrm{max}}=\infty,\quad\bm{\alpha}_{\mathrm{min}}=\epsilon,\quad\bm{\alpha}_{\mathrm{max}}=\infty,$ (16)
$\bm{\beta}_{\mathrm{min}}$ $=\epsilon,\quad\bm{\beta}_{\mathrm{max}}=\infty,\quad\bm{\gamma}_{\mathrm{min}}=\epsilon,\quad\bm{\gamma}_{\mathrm{max}}=\infty,$ (17)
@@ IV-C Box Constraints
We set the lower/upper bounds for $\bm{\bar{\kappa}}$ and $\mathbf{\bar{m}}$ as
$\bm{\bar{\kappa}}_{\mathrm{min}}$ $=\bm{\bar{\kappa}}_{0}-\mu\mathbf{e}_{\mathrm{be}},\quad\bm{\bar{\kappa}}_{\mathrm{max}}=\bm{\bar{\kappa}}_{0}+\mu\mathbf{e}_{\mathrm{be}},$ (18)
@@
where $\bm{\bar{\kappa}}_{0}$ and $\mathbf{\bar{m}}_{0}$ denote the initial rest curvature and rest twist (before parameter optimization), respectively, $\mu$ and $\eta$ the allowed change for rest curvature and rest twist, respectively, and $\mathbf{e}_{\mathrm{be}}$ and $\mathbf{e}_{\mathrm{tw}}$ vectors of all ones with the same dimensions as $\bm{\bar{\kappa}}$ and $\mathbf{\bar{m}}$, respectively. The stable range of $\bm{\bar{\kappa}}$ and thus $\mu$ can vary depending on simulation settings (e.g., strand geometry and material stiffness) [5]; we use $\mu=1$ by default based on our experiments. Additionally, since changes in rest twist are typically much smaller than those in rest curvature, we set $\eta=\mu/4$ to reduce the number of tunable parameters.
IV-D Newton-based Box-Constrained Optimization
We solve the primal, box-constrained nonlinear minimization subproblem (12) using a Newton-type optimizer. One could approximate the box constraints with a penalty objective to ensure fully unconstrained inner linear systems [15, 5], but penalty methods often require tedious parameter tuning to achieve acceptable results. Even with carefully tuned parameters, these methods may lead to compromises that violate box constraints. Therefore, instead of relying on penalty methods, we incorporate box constraints directly when computing the Newton directions, which is equivalent to solving box-constrained quadratic programs (BCQPs) [42].

### v2

@@ IV-D Twisting
$E_{\mathrm{tw},i}(\mathbf{y}_{i})=\frac{1}{2}\left(\frac{s\boldsymbol{\gamma}_{i}\pi r^{4}}{(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)(\mathbf{m}_{i}-\mathbf{\bar{m}}_{i})^{2},$ (4)
with $\boldsymbol{\gamma}_{i}$ representing the twisting stiffness parameter, and $\mathbf{m}_{i}$ and $\mathbf{\bar{m}}_{i}$ the twist (including the reference twist [7, 49, 46]) and rest twist, respectively.
V Parameter Optimization
Our goal is, at the initialization stage, to find a set of optimal rest shape and material stiffness parameters which achieve static equilibrium of the strands. Letting $\mathbf{p}$ be the optimization variable (to be detailed in Sec. V-A), our parameter optimization task is formulated as a constrained minimization problem:
@@ V Parameter Optimization
where $\mathbf{p}_{\mathrm{min}}$ and $\mathbf{p}_{\mathrm{max}}$ denote the box constraints (lower and upper bounds for $\mathbf{p}$, respectively) to ensure compliance with physical laws and limit significant parameter changes [5], and $\mathbf{c}(\mathbf{p})$ (defined in Sec. V-C) is a nonlinear hard constraint to enforce zero net forces. We define $\mathbf{p}_{0}$ as the initial value for $\mathbf{p}$ before optimization, and $R(\mathbf{p})$ acts as a regularizer to minimize the deviation of $\mathbf{p}$ from $\mathbf{p}_{0}$ with a diagonal weight matrix $\mathbf{W}$. We aim to solve (5) using ALM and a Newton-type optimizer (see Algorithm 1 in Sec. V-F for the procedure).
V-A Optimization Variable
We simultaneously optimize parameters of the rest shape ($\mathbf{\bar{l}},\boldsymbol{\bar{\kappa}},\mathbf{\bar{m}}$) and material stiffness ($\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}$) to achieve static equilibrium of an elastic strand. To improve robustness and efficiency, we use the following approaches to define the optimization variable $\mathbf{p}$.
Variable Scaling. While one typically specifies stiffness coefficients directly as $\mathbf{c}_{\mathrm{st}}(=s\boldsymbol{\alpha})$, $\mathbf{c}_{\mathrm{be}}(=s\boldsymbol{\beta})$, and $\mathbf{c}_{\mathrm{tw}}(=s\boldsymbol{\gamma})$, we instead optimize $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, and $\boldsymbol{\gamma}$. This approach addresses the significant scale differences between the stiffness coefficients ($\mathbf{c}_{\mathrm{st}}$, $\mathbf{c}_{\mathrm{be}}$, $\mathbf{c}_{\mathrm{tw}}$ on the order of $10^{9}$ [7, 46]) and rest shape parameters ($\mathbf{\bar{l}},\boldsymbol{\bar{\kappa}},\mathbf{\bar{m}}$ up to 10), enabling stable convergence of numerical optimizers and mitigating numerical issues. To ensure similar value scales, we initialize the stiffness scaling constant $s$ as $s=\frac{1}{3(N-2)}\sum_{i=1}^{N-2}(\mathbf{c}_{\mathrm{st},i}+\mathbf{c}_{\mathrm{be},i}+\mathbf{c}_{\mathrm{tw},i})$, and subsequently set $\boldsymbol{\alpha}=\mathbf{c}_{\mathrm{st}}/s$, $\boldsymbol{\beta}=\mathbf{c}_{\mathrm{be}}/s$, and $\boldsymbol{\gamma}=\mathbf{c}_{\mathrm{tw}}/s$. While equivalent formulations could be derived through proper scaling of (5) and (6), our change of variables ensures consistent scales in the Jacobians (see the supplementary material), thus avoiding catastrophic cancellation and enhancing numerical stability. In practice, we found that parameter optimization without this change of variables fails to achieve static equilibrium.
Variable Interleaving. Exploiting the one-dimensional structure and locally defined DER objectives of strands, we interleave the rest shape and stiffness parameters to ensure banded inner linear systems (which can be solved with fewer fill-ins via Cholesky-based direct solvers) within Newton-type optimizers, similar to the approach used in forward simulations [7]. Note that we exclude $\mathbf{\bar{l}}_{0}$ on the fixed edge from the parameter optimization [5].
Reduced Rest Curvatures. Furthermore, we optimize only two rest curvature DOFs per vertex, $\boldsymbol{\bar{\kappa}}_{i,0}$ and $\boldsymbol{\bar{\kappa}}_{i,1}$, while excluding $\boldsymbol{\bar{\kappa}}_{i,2}$ and $\boldsymbol{\bar{\kappa}}_{i,3}$, in contrast to previous work [5]. This choice stems from the association of $\boldsymbol{\bar{\kappa}}_{i,0}$ and $\boldsymbol{\bar{\kappa}}_{i,2}$ with the second material frame, while $\boldsymbol{\bar{\kappa}}_{i,1}$ and $\boldsymbol{\bar{\kappa}}_{i,3}$ relate to the first [6]. Their overlapping force spaces lead to similar final values for $\boldsymbol{\bar{\kappa}}_{i,0}$ and $\boldsymbol{\bar{\kappa}}_{i,2}$ without expanding the force spaces necessary to achieve static equilibrium, while including $\boldsymbol{\bar{\kappa}}_{i,2}$ and $\boldsymbol{\bar{\kappa}}_{i,3}$ would increase the optimization cost due to the additional DOFs. In practice, we synchronize these variables by setting $\boldsymbol{\bar{\kappa}}_{i,2}=\boldsymbol{\bar{\kappa}}_{i,0}$ and $\boldsymbol{\bar{\kappa}}_{i,3}=\boldsymbol{\bar{\kappa}}_{i,1}$ whenever $\boldsymbol{\bar{\kappa}}_{i,0}$ and $\boldsymbol{\bar{\kappa}}_{i,1}$ are updated. This approach effectively reduces the dimensionality of the rest curvature from 4D to 2D; however, we still use 4D rest curvatures in (3) to ensure the same dimensionality for curvature $\boldsymbol{\kappa}$ and rest curvature $\boldsymbol{\bar{\kappa}}$.
Based on these choices, we define the optimization variable for the parameters as $\mathbf{p}=(\boldsymbol{\bar{\kappa}}_{1,0},\boldsymbol{\bar{\kappa}}_{1,1},{\boldsymbol{\beta}}_{1},\mathbf{\bar{m}}_{1},{\boldsymbol{\gamma}}_{1},\mathbf{\bar{l}}_{1},{\boldsymbol{\alpha}}_{1},\ldots,\\ \boldsymbol{\bar{\kappa}}_{N-2,0},\boldsymbol{\bar{\kappa}}_{N-2,1},{\boldsymbol{\beta}}_{N-2},\mathbf{\bar{m}}_{N-2},{\boldsymbol{\gamma}}_{N-2},\mathbf{\bar{l}}_{N-2},{\boldsymbol{\alpha}}_{N-2})^{T}\in\mathbb{R}^{7N-14}$. The subset corresponding to the rest shape parameters consists of 4N84N-8 DOFs, which matches the number of active DOFs for the generalized positions $\mathbf{q}$. Any set of rest shape parameter values can exactly be converted into any corresponding set of generalized positions (and vice versa), and thus either representation could (potentially) be optimized for. We choose to use the rest shape parameters in the optimization because this approach offers more precise control over the magnitude of forces via the box constraints, and makes the Jacobians of the generalized forces with respect to the rest shape parameters much simpler [5].
V-B Augmented Lagrangian Method
@@
To handle the nonlinear constraints $\mathbf{c}(\mathbf{p})=0$ in (5), we use ALM [15, 36] and define an augmented Lagrangian objective with a Lagrange multiplier $\boldsymbol{\lambda}$ and a penalty parameter $\rho\ (>0)$:
@@
$L(\mathbf{p},\boldsymbol{\lambda})=R(\mathbf{p})-\boldsymbol{\lambda}^{T}\mathbf{c}(\mathbf{p})+\frac{\rho}{2}\left\lVert\mathbf{c}(\mathbf{p})\right\rVert_{2}^{2}.$ (6)
@@
$\mathbf{p},\boldsymbol{\lambda}=\operatorname*{arg\,min}_{\mathbf{p}_{\mathrm{min}}\leq\mathbf{p}\leq\mathbf{p}_{\mathrm{max}}}\operatorname*{arg\,max}_{\boldsymbol{\lambda}}L(\mathbf{p},\boldsymbol{\lambda}),$ (7)
which we optimize by iteratively solving primal and dual subproblems [15]. The primal minimization subproblem, with fixed $\boldsymbol{\lambda}^{k}$ for iteration index $k$ (initialized with $\boldsymbol{\lambda}^{0}=0$), is defined as
$\mathbf{p}^{k+1}=\operatorname*{arg\,min}_{\mathbf{p}_{\mathrm{min}}\leq\mathbf{p}\leq\mathbf{p}_{\mathrm{max}}}L(\mathbf{p},\boldsymbol{\lambda}^{k}),$ (8)
while the dual maximization subproblem can be solved via a simple vector update:
$\boldsymbol{\lambda}^{k+1}=\boldsymbol{\lambda}^{k}-\rho\mathbf{c}(\mathbf{p}^{k+1}).$ (9)
@@
V-C Zero Net Force Constraints
Given the total force due to the DER objectives, represented as $\mathbf{f}(\mathbf{p})=-\nabla E(\mathbf{q})$, it seems natural to define $\mathbf{c}(\mathbf{p})=\mathbf{f}(\mathbf{p})$ to enforce zero net forces. However, in this case, the term $\frac{\rho}{2}\left\lVert\mathbf{c}(\mathbf{p})\right\rVert_{2}^{2}$ in (6) becomes equivalent to a quadratic penalty on the total force, which is numerically ill-conditioned and fails to correctly account for the force contributions within the system [5]. To avoid these issues, we penalize the generalized forces in the kinetic energy metric [5] and define $\mathbf{c}(\mathbf{p})$ as
@@ V-C Zero Net Force Constraints
Consequently, $L(\mathbf{p},\boldsymbol{\lambda})$ in (6) can be rewritten as
$L(\mathbf{p},\boldsymbol{\lambda})=R(\mathbf{p})-\boldsymbol{\lambda}^{T}\mathbf{M}^{-\frac{1}{2}}\mathbf{f}(\mathbf{p})+\frac{\rho}{2}\left\lVert\mathbf{f}(\mathbf{p})\right\rVert^{2}_{\mathbf{M}^{-1}}.$ (11)
This objective (11) is numerically equivalent to the formulation proposed by Takahashi and Batty [5] when considering only the rest shape parameters (excluding stiffness parameters) and setting $\boldsymbol{\lambda}=0$. Thus, their formulation, which treats $\mathbf{c}(\mathbf{p})$ as a soft constraint, can be regarded as a subset of ours. In the following, we assume $\Delta t=1$ to simplify the formulations without affecting the optimization results [5].
V-D Box Constraints
We impose box constraints on the optimization parameters to prevent stability problems (in forward simulation) which can be caused by too significant changes in the parameter values. Specifically, we define $\mathbf{\bar{l}}_{\mathrm{min}}\leq\mathbf{\bar{l}}\leq\mathbf{\bar{l}}_{\mathrm{max}}$ (except for $\mathbf{\bar{l}}_{0}$), $\boldsymbol{\bar{\kappa}}_{\mathrm{min}}\leq\boldsymbol{\bar{\kappa}}\leq\boldsymbol{\bar{\kappa}}_{\mathrm{max}}$, $\mathbf{\bar{m}}_{\mathrm{min}}\leq\mathbf{\bar{m}}\leq\mathbf{\bar{m}}_{\mathrm{max}}$, $\boldsymbol{\alpha}_{\mathrm{min}}\leq\boldsymbol{\alpha}\leq\boldsymbol{\alpha}_{\mathrm{max}}$, $\boldsymbol{\beta}_{\mathrm{min}}\leq\boldsymbol{\beta}\leq\boldsymbol{\beta}_{\mathrm{max}}$, and $\boldsymbol{\gamma}_{\mathrm{min}}\leq\boldsymbol{\gamma}\leq\boldsymbol{\gamma}_{\mathrm{max}}$. Unlike the penalty-based approach [5], which requires safety margins to mitigate the risk of violating physical laws (e.g., negative rest lengths and material parameters), we handle box constraints precisely using MPRGP [16]. Consequently, we set lower and upper bounds for $\mathbf{\bar{l}}$, $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, and $\boldsymbol{\gamma}$ as
$\mathbf{\bar{l}}_{\mathrm{min}}$ $=\epsilon,\quad\mathbf{\bar{l}}_{\mathrm{max}}=\infty,\quad\boldsymbol{\alpha}_{\mathrm{min}}=\epsilon,\quad\boldsymbol{\alpha}_{\mathrm{max}}=\infty,$ (12)
$\boldsymbol{\beta}_{\mathrm{min}}$ $=\epsilon,\quad\boldsymbol{\beta}_{\mathrm{max}}=\infty,\quad\boldsymbol{\gamma}_{\mathrm{min}}=\epsilon,\quad\boldsymbol{\gamma}_{\mathrm{max}}=\infty,$ (13)
@@ V-D Box Constraints
We set the lower/upper bounds for $\boldsymbol{\bar{\kappa}}$ and $\mathbf{\bar{m}}$ as
$\boldsymbol{\bar{\kappa}}_{\mathrm{min}}$ $=\boldsymbol{\bar{\kappa}}_{0}-\mu\mathbf{e}_{\mathrm{be}},\quad\boldsymbol{\bar{\kappa}}_{\mathrm{max}}=\boldsymbol{\bar{\kappa}}_{0}+\mu\mathbf{e}_{\mathrm{be}},$ (14)
@@
where $\boldsymbol{\bar{\kappa}}_{0}$ and $\mathbf{\bar{m}}_{0}$ denote the initial rest curvature and rest twist (before parameter optimization), respectively, $\mu$ and $\eta$ the allowed change for rest curvature and rest twist, respectively, and $\mathbf{e}_{\mathrm{be}}$ and $\mathbf{e}_{\mathrm{tw}}$ vectors of all ones with the same dimensions as $\boldsymbol{\bar{\kappa}}$ and $\mathbf{\bar{m}}$, respectively. The stable range of $\boldsymbol{\bar{\kappa}}$ (i.e., which does not cause stability problems in forward simulations) and thus $\mu$ can vary depending on simulation settings (e.g., strand geometry and material stiffness) [5]; we use $\mu=1$ by default based on our experiments. Additionally, since changes in rest twist are typically much smaller than those in the rest curvature, we set $\eta=\mu/4$ to reduce the number of tunable parameters.
V-E Newton-based Box-Constrained Optimization
We solve the primal, box-constrained nonlinear minimization subproblem (8) using a Newton-type optimizer. While one could approximate the box constraints with a penalty objective to ensure fully unconstrained inner linear systems [5], penalty methods often require tedious parameter tuning to achieve acceptable results [15]. Even with carefully tuned parameters, these methods may lead to compromises that violate box constraints. Therefore, instead of relying on penalty methods, we incorporate box constraints directly when computing the Newton directions, which is equivalent to solving box-constrained quadratic programs (BCQPs) [50].